OpenStudy (anonymous):

A square has side length of x centimeters. A new rectangle is created by increasing one dimension by 4 centimeters and the other by 5 centimeters. What is the expanded and factored equations for the problem?

OpenStudy (anonymous):

someone please answer I really need help!

OpenStudy (anonymous):

equation for what?area or perimeter?

OpenStudy (anonymous):

area

OpenStudy (anonymous):

right area of rectangle=L*W L=x+5 W=x+4 area=(x+5)(x+4) =x^2+9x+20

OpenStudy (anonymous):

thank you so much